The speed of a car is decreased uniformly from 30. meters per second to 10. meters per second in 4.0 seconds. The magnitude of the car's acceleration is


1
Expert's answer
2021-02-24T12:49:47-0500
"a=\dfrac{|v_f-v_i|}{t}=\dfrac{|10\ \dfrac{m}{s}-30\ \dfrac{m}{s}|}{4\ s}=5\ \dfrac{m}{s^2}."